OpenStudy (anonymous):

How would I find the domain and range of (2)/(x^2-2x-3)? Is the domain x=-1 x=3

OpenStudy (cwrw238):

try factoring the denominator what values of x are not possible?

OpenStudy (anonymous):

x=-1 and x=3

OpenStudy (cwrw238):

the domain is all possible values of x

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

so -1 and 3 are not in the domain

OpenStudy (mathstudent55):

The domain is the values that x can have. y = 2/(x^2 - 2x - 3) has a denominator. Since division by zero is not allowed, you need to find which values of x would make the denominator zero. To do that, solve (which you did) x^2 - 2x - 3 = 0 (x - 3)(x + 1) = 0 x = 3 or x = -1 These are the value of x that must be excluded from the domain, so the domain is all real numbers except x = 3 and x = -1.

OpenStudy (anonymous):

oh so all reals except x=-1 and x=3 for the domain

OpenStudy (cwrw238):

yup

OpenStudy (anonymous):

ok thanks! but what about the range?

OpenStudy (cwrw238):

the range is all possible values of f(x)

OpenStudy (cwrw238):

so can you figure that out?

OpenStudy (anonymous):

um how would i set it up?

OpenStudy (anonymous):

dang ok.. don't solve it for me but how would i go about that?

OpenStudy (cwrw238):

well for a start the values of f(x) corresponding to x=3 or -1 will be excluded

OpenStudy (anonymous):

Couldn't I find the inverse of the function and then find the domain?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

yes the inverse of the function has domain which equals range of the function

OpenStudy (anonymous):

ok thanks for your time and patience!

OpenStudy (cwrw238):

yw